# Continuous GNN-based Anomaly Detection on Edge using Efficient Adaptive Knowledge Graph Learning

### v1

@@ II-A Video Anomaly Detection
State-of-the-art (SOTA) VAD models leverage advanced architectures like CNNs, LSTMs, and transformers to capture complex spatial-temporal patterns, enhancing anomaly detection accuracy [9, 22]. Recent approaches also integrate LLMs and vision models (LVMs) for improved contextual reasoning [27, 28]. However, LLM-based models are computationally expensive, limiting real-time applicability. MissionGNN addresses this by using a graph neural network (GNN) built on an LLM-generated knowledge graph, reducing computational costs while maintaining strong performance in real-time scenarios [26]. Yet, none of these models effectively handle continuous adaptation to evolving anomaly trends.
@@ II-B Edge Computing on VAD
Deploying VAD models on edge devices is critical for real-time surveillance in resource-constrained environments. Edge computing minimizes latency and dependence on cloud infrastructure, wich is essential for timely anomaly detection [16, 18]. While lightweight models enable real-time processing, they often underperform compared to SOTA models [23, 5]. MissionGNN offers a solution by encoding knowledge from large models into a lightweight graph neural network (GNN), reducing resource demands and enabling efficient edge deployment [26]. However, MissionGNN still requires cloud access to update its knowledge graph when anomalies change, limiting its utility in disconnected environments. Our approach allows continuous knowledge graph updates directly on edge devices, ensuring models remain effective without cloud dependency.
@@ III-C Light-weight GNN-based Decision Model Training
Our method employs hierarchical GNNs to reason over multiple knowledge graphs using a single frame. To extract richer video information, we use a transformer-based temporal model $\mathcal{T}\colon\mathbb{R}^{T\times D}\to\mathbb{R}^{D}$, processing sequences of embeddings from previous $T-1$ consecutive frames $\vec{F}_{t}=\{\vec{f}_{t-T+1},\vec{f}_{t-T+2},\cdots,\vec{f}_{t}\}\in\mathbb{R}^{T\times D}$ to the transformer and only takes the last output embedding $\vec{f}^{\prime}_{t}=\mathcal{T}(\vec{f}_{t})\in\mathbb{R}^{D}$ which is corresponding to the last input $\vec{f}_{t}$. Focusing on short-term relationships, the model assumes anomalies are detectable in brief intervals. The output of the short-term temporal model $\vec{f}^{\prime}_{t}$ is then processed by a decision model $f^{dec}\colon\mathbb{R}^{D}\to\mathbb{R}^{n+1}$ which is consists of a simple linear layer and softmax function:
@@ Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco.

### v2

@@ II-A Video Anomaly Detection
State-of-the-art (SOTA) VAD models leverage advanced architectures like CNNs, LSTMs, and transformers to capture complex spatial-temporal patterns, enhancing anomaly detection accuracy [10, 12]. Recent approaches also integrate LLMs and vision models (LVMs) for improved contextual reasoning [14, 19, 20, 21]. However, LLM-based models are computationally expensive, limiting real-time applicability. MissionGNN addresses this by using a graph neural network (GNN) built on an LLM-generated knowledge graph, reducing computational costs while maintaining strong performance in real-time scenarios [16]. Yet, none of these models effectively handle continuous adaptation to evolving anomaly trends.
@@ II-B Edge Computing on VAD
Deploying VAD models on edge devices is critical for real-time surveillance in resource-constrained environments. Edge computing minimizes latency and dependence on cloud infrastructure, wich is essential for timely anomaly detection [22, 23, 24, 25]. While lightweight models enable real-time processing, they often underperform compared to SOTA models [26, 27]. MissionGNN offers a solution by encoding knowledge from large models into a lightweight graph neural network (GNN), reducing resource demands and enabling efficient edge deployment [16]. However, MissionGNN still requires cloud access to update its knowledge graph when anomalies change, limiting its utility in disconnected environments. Our approach allows continuous knowledge graph updates directly on edge devices, ensuring models remain effective without cloud dependency.
@@ III-C Light-weight GNN-based Decision Model Training
Our method employs hierarchical GNNs to reason over multiple knowledge graphs using a single frame. To extract richer video information, we use a transformer-based temporal model $\mathcal{T}\colon\mathbb{R}^{T\times D}\to\mathbb{R}^{D}$, processing sequences of embeddings from previous $T-1$ consecutive frames $\vec{F}_{t}=\{\vec{f}_{t-T+1},\vec{f}_{t-T+2},\cdots,\vec{f}_{t}\}\in\mathbb{R}^{T\times D}$ to the transformer and only takes the last output embedding $\vec{f}^{\prime}_{t}=\mathcal{T}(\vec{f}_{t})\in\mathbb{R}^{D}$ which is corresponding to the last input $\vec{f}_{t}$. Focusing on short-term relationships, the model assumes anomalies are detectable in brief intervals.
The output of the short-term temporal model $\vec{f}^{\prime}_{t}$ is then processed by a decision model $f^{dec}\colon\mathbb{R}^{D}\to\mathbb{R}^{n+1}$ which is consists of a simple linear layer and softmax function:
@@ Acknowledgements
This work was supported in part by the DARPA Young Faculty Award, the National Science Foundation (NSF) under Grants #2127780, #2319198, #2321840, #2312517, and #2235472, the Semiconductor Research Corporation (SRC), the Office of Naval Research through the Young Investigator Program Award, and Grants #N00014-21-1-2225 and #N00014-22-1-2067, Army Research Office Grant #W911NF2410360, and the U.S. Army Combat Capabilities Development Command (DEVCOM) Army Research Laboratory under Support Agreement No. USMA 21050. Additionally, support was provided by the Air Force Office of Scientific Research under Award #FA9550-22-1-0253, along with generous gifts from Xilinx and Cisco. The views expressed in this paper are those of the authors and do not reflect the official policy or position of the U.S. Military Academy, the U.S. Army, the U.S. Department of Defense, or the U.S. Government.